OpenStudy (anonymous):

There are 8 people on the debate team. In how many ways can the coach choose 4 members to send to competition? 70 1,680 40,320

OpenStudy (anonymous):

8! 8*7*6*5*4*3*2*1 = 40320

OpenStudy (anonymous):

wait nevermind thats wrong

OpenStudy (anonymous):

@somuchhomework that counts how many ways he can arrange a line of the 8 students.

OpenStudy (anonymous):

It'd be \(\dbinom83=\dfrac{8!}{3!(8-3)!}\)

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

i got 112 for that though tf

OpenStudy (campbell_st):

but aren't you selecting 4 people for the team... and not 3

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh sorry, @campbell_st is right, it should be \[\binom84=\frac{8!}{4!(8-4)!}\]

OpenStudy (anonymous):

how do i get the answer from that?